A Professor of Economics from the Department of Economics Education of the Faculty of Social Sciences Education of the University of Education, Winneba, has estimated the demand and supply functions for students’ accommodation on campus to be respectively:

10P + 3Q = 60 and

P = Q – 0.5

Further, the professor claims that the average cost of Hostel operators is




1
Expert's answer
2021-07-26T07:13:02-0400

Demand function

10P+3Q=60

Supply function

P=Q-0.5

=10(Q-0.5)+3Q=60

13Q-5=60

13Q=60+5

13Q=65

Q=5

P=5-0.5=4.5